# Improving Deep Reinforcement Learning by Reducing the Chain Effect of Value and Policy Churn

### v1

@@ 1 Introduction
One fundamental recipe for the success of Deep Reinforcement Learning (DRL) is powerful approximation and generalization provided by deep neural networks, which augments the ability of RL with tabular or linear approximation to large state spaces. However, on the other side of this benefit is less control over the function dynamics. Network outputs can change indirectly to unexpected values after any random batch update for input data not included in the batch, called churn in this paper. This change is particularly problematic for an RL agent due to its non-stationary nature, which can exacerbate instability, suboptimality and even collapse. Therefore, it is of significance to understand and control these undesired dynamics to address learning issues and improve performance.
@@
Consistent efforts have been devoted by RL community to gain better understandings of the learning dynamics from different perspectives (Achiam et al., 2019; Kumar et al., 2022; Liu et al., 2023; Lyle et al., 2022b). Recently, Schaul et al. (2022) studied a novel churn phenomenon in the learning process of typical value-based RL algorithms like DoubleDQN (van Hasselt et al., 2016). The phenomenon reveals that the greedy actions of about $10\%$ of states in the replay buffer change after a single regular batch update. Such a dramatic churn can persist throughout the learning process of DoubleDQN, causing instabilities.
@@
In this paper, we aim to take a step further to understand how churn occurs and influences learning in different DRL settings beyond value-based RL, as well as to propose a method to control churn. We start by characterizing churn formally in a view of Generalized Policy Iteration (GPI) with function approximation, to best cover most DRL settings. The impact of churn is two-fold in this view: the churn in policy improvement (called the policy churn) changes policy outputs on states that are not directly updated; while the churn in value estimation (called the value churn) also changes action-value landscape, thus altering greedy action and action gradient. We then discover a chain effect of churn that exhibits a cycle where the two types of churn compound and bias the learning dynamics throughout the iteration. Further, we move on from the general analysis to concrete DRL settings. We focus on the learning issues caused by the chain effect including greedy action deviation in value-based methods, trust region violation in proximal policy optimization (Schulman et al., 2017) and dual bias of policy value in actor-critic methods. The connection between the chain effect of churn and the issues necessitates an explicit control of the churn.
@@
To this end, we propose a method called Churn Approximated ReductIoN (CHAIN), to reduce the chain effect of churn across different DRL settings. The main idea of CHAIN is to reduce the undesirable changes to the outputs of the policy and value networks for states (and actions) outside of the current batch of data for regular DRL training. This reduction is achieved by minimizing the change in target values for a separate batch of data when optimizing the original policy or value learning objective. CHAIN is easy to implement and plug in most existing DRL algorithms with only a few lines of code. In our experiments, we evaluate the efficacy of CHAIN in a range of environments, including MinAtar (Young and Tian, 2019), OpenAI MuJoCo (Brockman et al., 2016), DeepMind Control Suite (Tassa et al., 2018) and D4RL (Fu et al., 2020). The results show that our method can effectively reduce churn and mitigate the learning issues, thus improving sample efficiency or final performance across online and offline, value-based and policy-based DRL settings. Moreover, our results also show that our method helps to scale DRL agents up and achieves significantly better learning performance when using wider or deeper networks.
@@
The main contributions of this work are summarized as follows: (1) We study how churn occurs and influences learning from the perspective of GPI with function approximation and present the chain effect of churn. (2) We show how churn results in three types of learning issues in typical DRL settings. (3) We propose a simple and general method and demonstrate its effectiveness in reducing churn and improving learning performance across a range of DRL settings and environments.
@@ 2 Prior Work
Among these directions, one notable thread is to understand the learning dynamics of DRL agents with a focus on the non-stationary nature of RL. A prominent phenomenon of representation ability loss is studied in (Dabney et al., 2021; Igl et al., 2021; Kumar et al., 2021, 2022; Ma et al., 2023), which reveals how representations become less useful in later stages of learning, leading to myopic convergence. Further, empirical studies in (Nikishin et al., 2022; D’Oro et al., 2023; Sokar et al., 2023; Nauman et al., 2024) demonstrate that the loss of approximation ability becomes severe and leads to collapse when high replay-ratios are adopted for better sample efficiency, while network resets and normalization methods can be simple and effective remedies. This is further identified as plasticity loss in DRL (Lyle et al., 2022a; Abbas et al., 2023; Dohare et al., 2023; Lyle et al., 2024; Xu et al., 2024).
Recently, it has been found that there is a dramatic change in the policy distribution where a large portion of the greedy actions change after each batch update, called policy churn Schaul et al. (2022). Meanwhile, it also presents a lack of understanding of churn’s effect on the learning behaviors of DRL agents regarding stability, convergence, exploration, etc. Later, Kapturowski et al. (2023) takes the inspiration and proposes a method to robustify the agent’s behavior by adding an additional policy head to the value network that fits the $\epsilon$-greedy policy via policy distillation. In this work, we further the study of churn in a more general formal framework, where the churn occurs in both value and policy learning. In particular, we focus on the dynamics of churn during the learning process and how it incurs issues in different DRL algorithms and propose a practical method to reduce churn and improve learning performance.
@@ 4.1 Generalized Policy Iteration under Churn
Generalized Policy Iteration (GPI) (Sutton and Barto, 1988) is widely used to refer to the general principle of learning in an Evaluation-Improvement iteration manner, which applies to almost all RL methods. In the context of DRL, i.e., with network representation and mini-batch training, the value and policy networks’ outputs can have unexpected changes, i.e., the churn, after each mini-batch training for the states not included in the batch. Such churns are neglected in most DRL methods, let alone their influence on the practical learning process. In Figure 1, we extend the classic GPI diagram by taking churn into consideration, to show how it is involved in the learning process intuitively.
In the evaluation process, the parameterized $Q$-network $Q_{\theta}$ approximates the value of the current policy via repeated mini-batch training. Under the impact of churn, the $Q$-network is not likely to have output predictions exactly as what was updated with explicit mini-batch training. For example, let’s imagine we have a virtual network $\bar{Q}_{\theta}$ that only accepts the changes for the states updated by mini-batch training directly and remains unchanged for the others.
@@
Thus, the value churn, denoted by $\bar{Q}_{\theta}\leadsto Q_{\theta}$, is an implicit process that alters the virtual network $\bar{Q}_{\theta}$ to the approximation $Q_{\theta}$ we obtained in practice. Similarly, the policy churn $\bar{\pi}_{\phi}\leadsto{\pi}_{\phi}$ occurs in the improvement process. As illustrated in Figure 1, the value churn and the policy churn are interwoven in the Evaluation-Improvement process, and usually, we can assume that the churns make non-negligible changes, i.e., $Q_{\theta}\neq\bar{Q}_{\theta}$ and $\pi_{\phi}\neq\bar{\pi}_{\phi}$. Therefore, we delve into the cause of churn, its impact on learning, and possible remedies to mitigate its negative impact in the following sections.
@@ 4.2 Definition of Value and Policy Churn
Arguably, churn is an innate property of neural networks and it is closely related to problems like interference (Liu et al., 2020, 2023) and catastrophic forgetting (Lan et al., 2023) in different contexts.
@@
Now we characterize $\mathcal{C}_{Q}$ and $\mathcal{C}_{\pi}$ as functions of $\Delta_{\theta},\Delta_{\phi}$ with the help of Neural Tangent Kernel (NTK) expression (Achiam et al., 2019). For clarity, we use $B_{\text{train}}=\{s,a\}$ and $B_{\text{ref}}=\{\bar{s},\bar{a}\}$ and abbreviate $B_{\text{train}},B_{\text{ref}}$ and step size $\alpha$ when context is clear. Concretely,
@@
Eq. 2 shows that the value and policy churn is mainly determined by the kernels of the $Q$-network $k_{\theta}$ and the policy network $k_{\phi}$, along with the TD error and the action gradient. This indicates that churn is determined by both the network’s property itself and the learning we performed with the network.
@@ 4.3 From Single-step Interplay to The Chain Effect of Churn
From these aspects, we can deduce the difference between the parameter updates under the impact of the value and policy churn (denoted by $\tilde{\Delta}_{\theta},\tilde{\Delta}_{\phi}$) and the conventional ones $\Delta_{\theta},\Delta_{\phi}$. As a result, we can find that the value and policy churn, as well as the deviations derived, introduce biases in the parameter updates. We provide the complete discussion and derivation in Section A.2.
@@ 5.1 Consequences of the Chain Effect of Churn in Different DRL Scenarios
Since churn is involved in most DRL methods as illustrated by Figure 1, we concretize our study and focus on several typical DRL scenarios below.
@@
Trust region plays a critical role in many policy gradient methods for reliable and efficient policy updates. Proximal Policy Optimization (PPO) (Schulman et al., 2017) uses a clipping mechanism as a simple but effective surrogate of the trust region for TRPO (Schulman et al., 2015): Clip$\big{(}r(\phi_{\text{old}},\phi),1-\epsilon,1+\epsilon\big{)}$ and $r(\phi_{\text{old}},\phi)=\frac{\pi_{\phi}(a|s)}{\pi_{\phi_{\text{old}}}(a|s)}$. With respect to policy churn, even though the PPO policy conforms to the trust region for the states in the current training batch, it could silently violate the trust region for other states, including previously updated ones. Consider the policy update $\phi\rightarrow\phi^{\prime}$, it is highly likely to have $r(\phi,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})=\frac{\pi_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}}}(\bar{a}|\bar{s})}{\pi_{\phi}(\bar{a}|\bar{s})}\neq 1$ and thus $r(\phi_{\text{old}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})=\frac{\pi_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}}}(\bar{a}|\bar{s})}{\pi_{\phi_{\text{old}}}(\bar{a}|\bar{s})}=r(\phi_{\text{old}},\phi)r(\phi,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})\neq r(\phi_{\text{old}},\phi)$. Since we have no information about $r(\phi,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})$, there is no guarantee for the trust region $1-\epsilon\leq r(\phi_{\text{old}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})\leq 1+\epsilon$ to be respected after churn. Intuitively, this silent violation is hazardous and detrimental to learning.
@@ 5.2 A Regularization Method for Churn Reduction
By minimizing $L_{\text{QC}}$ and $L_{\text{PC}}$, we can reduce the value and policy churn and suppress the chain effect further. This allows us to use the churn reduction regularization terms along with standard RL objectives, and arrives at DRL with CHAIN finally:
@@
where $B_{\text{train}},B_{\text{ref}}$ are two separate batches randomly sampled from $D$, and $\lambda_{Q},\lambda_{\pi}$ are coefficients that control the degree of regularization. CHAIN serves as a plug-in component that can be implemented with only a few lines of code modification in most DRL methods. The pseudocodes are omitted here, and we refer the readers to Algorithm 1 in the Appendix.
@@
To alleviate the difficulty of manually selecting the regularization coefficients, we add a simple but effective method to adjust $\lambda_{Q},\lambda_{\pi}$ adaptively during the learning process. The key principle behind this is to keep a consistent relative scale (denoted by $\beta$) between the churn reduction regularization terms and the original DRL objectives. Concrete, by maintaining the running means of the absolute $Q$ loss $|\bar{L}_{Q}|$ and the VCR term $|\bar{L}_{\text{QC}}|$, $\lambda_{Q}$ is computed dynamically as $\lambda_{Q}=\beta\frac{|\bar{L}_{Q}|}{|\bar{L}_{\text{QC}}|}$, which is similar for $\lambda_{\pi}$. This is inspired by our empirical observations and the recent study on addressing the scale difference across different domains (Hafner et al., 2023).
@@ 6 Experiments
We organize our experiments into the four subsections below that correspond to the three DRL scenarios discussed in Section 5.1 as well as a DRL scaling setting. Our experiments include 20 online RL tasks from MinAtar, MuJoCo, DMC, and 8 offline RL datasets from D4RL, as well as 6 popular algorithms, i.e., DoubleDQN, PPO, TD3, SAC, IQL, AWAC. We provide the experimental details in Appendix B and more results in Appendix C.

### v2

@@ 1 Introduction
One fundamental recipe for the success of Deep Reinforcement Learning (DRL) is powerful approximation and generalization provided by deep neural networks, which augments the ability of RL with tabular or linear approximation to large state spaces. However, on the other side of this benefit is less control over the function dynamics. Network outputs can change indirectly to unexpected values after any random batch update for input data not included in the batch, called churn in this paper. This change is particularly problematic for an RL agent due to its non-stationary nature, which can exacerbate instability, suboptimality, and even collapse. Therefore, it is important to understand and control these undesired dynamics to address learning issues and improve performance.
@@
Consistent efforts have been devoted by the RL community to gain a better understanding of the learning dynamics from different perspectives (Achiam et al., 2019; Kumar et al., 2022; Liu et al., 2023; Lyle et al., 2022b). Recently, Schaul et al. (2022) studied a novel churn phenomenon in the learning process of typical value-based RL algorithms like DoubleDQN (van Hasselt et al., 2016). The phenomenon reveals that the greedy actions of about $10\%$ of states in the replay buffer change after a single regular batch update. Such a dramatic churn can persist throughout the learning process of DoubleDQN, causing instabilities.
@@
In this paper, we aim to take a step further to understand how churn occurs and influences learning in different DRL settings beyond value-based RL, as well as to propose a method to control churn. We start by formally characterizing churn in view of Generalized Policy Iteration (GPI) with function approximation to best cover most DRL settings. The impact of churn is two-fold in this view: the churn in policy improvement (called the policy churn) changes policy outputs on states that are not directly updated, while the churn in value estimation (called the value churn) also changes the action-value landscape, thus altering greedy action and action gradient. We then discover a chain effect of churn that exhibits a cycle where the two types of churn compound and bias the learning dynamics throughout the iteration. Further, we move on from the general analysis to concrete DRL settings. We focus on the learning issues caused by the chain effect including greedy action deviation in value-based methods, trust region violation in proximal policy optimization (Schulman et al., 2017) and dual bias of policy value in actor-critic methods. The connection between the chain effect of churn and the issues necessitates an explicit control of the churn.
@@
To this end, we propose a method called Churn Approximated ReductIoN (CHAIN) to reduce the chain effect of churn across different DRL settings. The main idea of CHAIN is to reduce the undesirable changes to the outputs of the policy and value networks for states (and actions) outside of the current batch of data for regular DRL training. This reduction is achieved by minimizing the change in target values for a separate batch of data when optimizing the original policy or value learning objective. CHAIN is easy to implement and plug in most existing DRL algorithms with only a few lines of code [^1] https://github.com/bluecontra/CHAIN. In our experiments, we evaluate the efficacy of CHAIN in a range of environments, including MinAtar (Young and Tian, 2019), OpenAI MuJoCo (Brockman et al., 2016), DeepMind Control Suite (Tassa et al., 2018) and D4RL (Fu et al., 2020). The results show that our method can effectively reduce churn and mitigate the learning issues, thus improving sample efficiency or final performance across online and offline, value-based and policy-based DRL settings. Moreover, our results also show that our method helps to scale DRL agents up and achieves significantly better learning performance when using wider or deeper networks.
@@
The main contributions of this work are summarized as follows: (1) We study how churn occurs and influences learning from the perspective of GPI with function approximation and present the chain effect of churn. (2) We show how churn results in three learning issues in typical DRL settings. (3) We propose a simple and general method and demonstrate its effectiveness in reducing churn and improving learning performance across various DRL settings and environments.
@@ 2 Prior Work
One notable thread is to understand the learning dynamics of DRL agents with a focus on the non-stationary nature of RL. A prominent phenomenon of representation ability loss is studied in (Dabney et al., 2021; Igl et al., 2021; Kumar et al., 2021, 2022; Ma et al., 2023), which reveals how representations become less useful in later stages of learning, leading to myopic convergence. Further, empirical studies in (Nikishin et al., 2022; D’Oro et al., 2023; Sokar et al., 2023; Nauman et al., 2024) demonstrate that the loss of approximation ability becomes severe and leads to collapse when high replay-ratios are adopted for better sample efficiency, while network resets and normalization methods can be simple and effective remedies. This is further identified as plasticity loss in DRL (Lyle et al., 2022a; Abbas et al., 2023; Dohare et al., 2023; Lyle et al., 2024; Xu et al., 2024).
Recently, it has been found that there is a dramatic change in the policy distribution where a large portion of the greedy actions change after each batch update, called policy churn Schaul et al. (2022). Although intuitively related to generalization (Bengio et al., 2020) and interference (Liu et al., 2023), it presents a lack of understanding of churn’s effect on the learning behaviors of DRL agents regarding stability, convergence, exploration, etc. Kapturowski et al. (2023) takes the inspiration and proposes a method to robustify the agent’s behavior by adding an additional policy head to the value network that fits the $\epsilon$-greedy policy via policy distillation. In this work, we further the study of churn in a more general formal framework, where churn occurs in both value and policy learning. In particular, we focus on the dynamics of churn during the learning process and how it incurs issues in different DRL algorithms and propose a practical method to reduce churn and improve learning performance.
@@ 4.1 Generalized Policy Iteration under Churn
Generalized Policy Iteration (GPI) (Sutton and Barto, 1988) is widely used to refer to the general principle of learning in an Evaluation-Improvement iteration manner, which applies to almost all RL methods. In the context of DRL, i.e., with network representation and mini-batch training, the value and policy networks’ outputs can have unexpected changes, i.e., the churn, after each mini-batch training for the states not included in the batch. Such churns are neglected in most DRL methods, let alone their influence on the practical learning process. In Figure 1, we extend the classic GPI diagram by considering churn to show how it is involved in the learning process intuitively.
In the evaluation process, the parameterized $Q$-network $Q_{\theta}$ approximates the value of the current policy via repeated mini-batch training. Under the impact of churn, the $Q$-network is not likely to have output predictions the same as what was updated with explicit mini-batch training. For example, let’s imagine we have a virtual network $\bar{Q}_{\theta}$ that only accepts the changes for the states updated by mini-batch training directly and remains unchanged for the others.
@@
Thus, the value churn, denoted by $\bar{Q}_{\theta}\leadsto Q_{\theta}$, is an implicit process that alters the virtual network $\bar{Q}_{\theta}$ to the approximation $Q_{\theta}$ we obtained in practice. Similarly, the policy churn $\bar{\pi}_{\phi}\leadsto{\pi}_{\phi}$ occurs in the improvement process. As illustrated in Figure 1, the value churn and the policy churn are interwoven in the Evaluation-Improvement process. Usually, we can assume that the churns make non-negligible changes, i.e., $Q_{\theta}\neq\bar{Q}_{\theta}$ and $\pi_{\phi}\neq\bar{\pi}_{\phi}$. Therefore, we delve into the cause of churn, its impact on learning, and possible remedies to mitigate its negative impact in the following sections.
@@ 4.2 Definition of Value and Policy Churn
Arguably, churn is an innate property of neural networks, and it is closely related to problems like interference (Liu et al., 2020, 2023) and catastrophic forgetting (Lan et al., 2023) in different contexts.
@@
Now we characterize $\mathcal{C}_{Q}$ and $\mathcal{C}_{\pi}$ as functions of $\Delta_{\theta},\Delta_{\phi}$ with the help of Neural Tangent Kernel (NTK) (Achiam et al., 2019). For clarity, we use $B_{\text{train}}=\{s,a\}$ and $B_{\text{ref}}=\{\bar{s},\bar{a}\}$ and abbreviate $B_{\text{train}},B_{\text{ref}}$ and step size $\alpha$ when context is clear. Concretely,
@@
Eq. 2 shows that the value and policy churn are mainly determined by the kernels of the $Q$-network $k_{\theta}$ and the policy network $k_{\phi}$, along with the TD error and the action gradient. This indicates that churn is determined by both the network’s property itself and the learning we performed with the network.
@@ 4.3 From Single-step Interplay to The Chain Effect of Churn
From these aspects, we can deduce the difference between the parameter updates under the impact of the value and policy churn (denoted by $\tilde{\Delta}_{\theta},\tilde{\Delta}_{\phi}$) and the conventional ones $\Delta_{\theta},\Delta_{\phi}$. As a result, we can find that the value and policy churn, as well as the deviations derived, introduce biases in the parameter updates. We provide the complete discussion and derivation in Section B.2.
@@ 5.1 Consequences of the Chain Effect of Churn in Different DRL Scenarios
Since churn is involved in most DRL methods, as illustrated by Figure 1, our study focuses on several typical DRL scenarios below.
@@
Trust region plays a critical role in many policy gradient methods for reliable and efficient policy updates. Proximal Policy Optimization (PPO) (Schulman et al., 2017) uses a clipping mechanism as a simple but effective surrogate of the trust region for TRPO (Schulman et al., 2015): Clip$\big{(}r(\phi_{\text{old}},\phi),1-\epsilon,1+\epsilon\big{)}$ and $r(\phi_{\text{old}},\phi)=\frac{\pi_{\phi}(a|s)}{\pi_{\phi_{\text{old}}}(a|s)}$. With respect to policy churn, even though the PPO policy conforms to the trust region for the states in the current training batch, it could silently violate the trust region for other states, including previously updated ones. Consider the policy update $\phi\rightarrow\phi^{\prime}$, it is highly likely to have $r(\phi,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})=\frac{\pi_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}}}(\bar{a}|\bar{s})}{\pi_{\phi}(\bar{a}|\bar{s})}\neq 1$ and thus $r(\phi_{\text{old}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})=\frac{\pi_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}}}(\bar{a}|\bar{s})}{\pi_{\phi_{\text{old}}}(\bar{a}|\bar{s})}=r(\phi_{\text{old}},\phi)r(\phi,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})\neq r(\phi_{\text{old}},\phi)$. Since we have no information about $r(\phi,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})$, there is no guarantee for the trust region $1-\epsilon\leq r(\phi_{\text{old}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\phi^{\prime}})\leq 1+\epsilon$ to be respected after churn. Intuitively, this implicit violation is hazardous and detrimental to learning.
@@ 5.2 A Regularization Method for Churn Reduction
By minimizing $L_{\text{QC}}$ and $L_{\text{PC}}$, we can reduce the value and policy churn and suppress the chain effect further. This allows us to use the churn reduction regularization terms along with standard RL objectives and arrives at DRL with CHAIN finally:
@@
where $B_{\text{train}},B_{\text{ref}}$ are two separate batches randomly sampled from $D$, and $\lambda_{Q},\lambda_{\pi}$ are coefficients that control the degree of regularization. CHAIN serves as a plug-in component that can be implemented with only a few lines of code modification in most DRL methods. The pseudocode is omitted here, and we refer the readers to Algorithm 1 in the Appendix.
@@
To alleviate the difficulty of manually selecting the regularization coefficients, we add a simple but effective method to adjust $\lambda_{Q},\lambda_{\pi}$ adaptively during the learning process. The key principle behind this is to keep a consistent relative scale (denoted by $\beta$) between the churn reduction regularization terms and the original DRL objectives. More precisely, by maintaining the running means of the absolute $Q$ loss $|\bar{L}_{Q}|$ and the VCR term $|\bar{L}_{\text{QC}}|$, $\lambda_{Q}$ is computed dynamically as $\lambda_{Q}=\beta\frac{|\bar{L}_{Q}|}{|\bar{L}_{\text{QC}}|}$, which is similar for $\lambda_{\pi}$. This is inspired by our empirical observations and the recent study on addressing the reward scale difference across different domains (Hafner et al., 2023).
@@ 6 Experiments
We organize our experiments into the four subsections below that correspond to the three DRL scenarios discussed in Section 5.1 as well as a DRL scaling setting. Our experiments include 20 online RL tasks from MinAtar, MuJoCo, DMC, and 8 offline RL datasets from D4RL, as well as 6 popular algorithms, i.e., DoubleDQN, PPO, TD3, SAC, IQL, AWAC. We provide the experimental details in Appendix C and more results in Appendix D.